# Vulnerability Disclosure or Notification? Best Practices for Reaching Stakeholders at Scale

### v1

@@ 4.1. Stages of Operation
In the literature selection, we cross-reviewed each selected work in detail and aimed to figure out the common procedure for performing vulnerability disclosure and notification. Eventually, we compiled five stages to represent the procedure implemented across the literature. We then use the stages to extract key points, numbers, and remarks from the selected work to better understand efforts, considerations and reviews made in the operations. This results in the following stages to assess the selected publications:
Pre-Assessment – Before a vulnerability disclosure or notification can start, a finder will assess the risk and impact of the discovered vulnerability. We identify the type and number of involved stakeholders and the vulnerabilities, and extract the impact scale of the vulnerability to understand the preparation before an operation.
@@
Post-Review – An operation can be reviewed by tracking the remediation rate, feedback of stakeholders, and experiences in each stage. A finder can also reflect on the operation and contribute to best practices. We extract the notified hosts and domains, remediation rates, interactions between the stakeholders, and contribution to best practices of large-scale disclosure and notification operations.
To present the overview of the assessment of each publication, we arrange three tables based on the above 5 stages. Tables 1 and 2 both contain the publication titles, pre-assessment, and post-review to indicate the efforts and outcomes from each work with different scenarios considered. Table 3 shows the single or multiple communication channels and messaging infrastructure implemented in each reviewed disclosure or notification operation.
@@ 5. Lessons Learned
This section presents the insights we have gained from the selected work on large-scale vulnerability disclosure and notification. We also integrate the experience we learn from the community, support organisations, existing guidelines and standards to present the essence of the disclosure or notification procedure.
With Table 1, 2, and 3 presented as an overview of disclosure and notification operations, we first discuss the development of best practices from the experiences of the selected work in Section 5.1. Then, we compare how the finders adopted the past best practices and what they learned from their operations in Section 5.2.
@@ 5.1. The Development of Best Practices
Security researchers, practitioners, communities, and support organisations have been seeking optimal methods and strategies based on the disclosure and notification best practices presented in their time, as shown in Table 1 and 2. We have seen changes in their selection of communication channels, messaging infrastructure, contact retrieval, message formulation, and disclosure policies depending on the nature of the found vulnerabilities, notification scalability, and recommendations from the communities.
@@
The essential part of vulnerability disclosure and notification is ensuring messages reach the responsible parties and prompt them to action. As presented in Table 3, the finders used different communication channels and messaging infrastructure to ensure the delivery of the disclosure messages, either on their own or with support organisations, such as national CSIRTs and Shadowserver Foundation. To maximise the notification coverage, all work adopted more than a single communication channel, and most of them delegated notifications to support organisations.
@@
Shadowserver Foundation is a non-profit organisation founded in 2004, driven by the vision of a secure, threat-free internet. They do this by scanning the internet for known vulnerabilities and performing notifications, managing sinkholes for malware command-and-control infrastructure, and operating honeypots and honeyclients to monitor threat developments. Over the years, the team at Shadowserver Foundation has created a large, trusted network with national CSIRTs, ISPs and others to ensure that they are able to reach as many stakeholders as possible. Anyone can sign up to get threat reports about their domain or IP as long as they can prove ownership. More importantly, Shadowserver Foundation can support researchers in performing notifications to stakeholders, as described by Beitis et al. (Beitis and Vanhoef, [n. d.]; The Shadowserver Foundation, [n. d.]).
The earlier vulnerability notification operations in (Durumeric et al., 2014; Li et al., 2016) focused on email and coordinator delegation notifications, then others in (Çetin et al., 2017a; Stock et al., 2018b, 2016a) extensively implemented diverse communication channels, compared the effectiveness of each empirically, and provided valuable suggestions to best practices on large-scale notifications. These operations served as the foundation for the large-scale notification best practices.
The later operations in (Chen et al., 2024; Stivala et al., 2024) adopted best practices from the prior work. They narrowed down the channel selection and dived into the experiences with email notifications. On the contrary, Maaß et al. (Maass et al., 2021) decided to explore alternatives other than the best practices with costly manual efforts on email, post, and phone call. All their results reflected the prior best practices and helped improve notification best practices at scale. In all the above work, locating accurate contact information remains the major challenge regardless of the selected communication channels. On the other hand, the researchers (Çetin et al., 2019b; Rodríguez et al., 2021) cooperated with a medium-sized ISP to explore the effectiveness of email and specialised ISP notification systems with different vulnerabilities, which serves as an example in notification with direct intervention from an ISP to their customers.
Furthermore, vulnerability disclosure to vendors has received promising best practices as implemented in  (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Boucher and Anderson, 2022; Van Hove et al., 2023). Boucher et al. (Boucher and Anderson, 2022) have provided well-structured details on acquiring CVEs, CERT/CC VINCE notification, public press disclosure, and constraints in academic publication. However, despite the adoption of best practices, the researchers in the two disclosure operations (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Van Hove et al., 2023) tackled different software vulnerabilities but encountered the challenge of inviting the same stakeholder to participate in CVD. Last but not least, the two operations (Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]) cover both vulnerability disclosure and notification. Their experiences revealed the long-lasting challenges of starting from disclosure to large-scale notification and attempting newer communication channels and strategies to mitigate the vulnerabilities in time.
5.2. The Comparison of Different Scenarios
To understand how the finders in the selected work adopted and contributed to best practices, we describe and compare their experiences based on the assessment stages in Section 4.1. The comparison follows the order of the five stages: Pre-Assessment, Communication Channel, Messaging Infrastructure, Disclosure Policy and Message, and Post-Review.
5.2.1. Pre-Assessment
All disclosure operations (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Boucher and Anderson, 2022; Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]) initiated the coordinated vulnerability disclosure directly to affected vendors with newly found vulnerabilities and later delegated their disclosure to support organisations. Before their disclosure, most finders could identify several vendors with bug bounty (Boucher and Anderson, 2022) or tracking program (Moura and Heidemann, 2023), and a short list of direct disclosure contact information (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Boucher and Anderson, 2022; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]; Moura and Heidemann, 2023), which ranges from 5 to 14 affected parties due to limited capacity or testing environments. However, most finders faced challenges such as limited contact information of other stakeholders and the potential development to large-scale disclosure. They then consulted national CSIRTs, CERT/CC (VINCE), or NCSC-NL, except for the researchers in (Moura and Heidemann, 2023). The impact scale later increased dramatically from the cooperation with support organisations, which is presented in the Post-Review in Table 1. Despite the lack of contact information in the pre-assessment, the disclosure parties could still mitigate the challenges with the support organisations.
The notification operations are vulnerability notifications to the stakeholders, with hosts or domains remaining vulnerable after public disclosure. The finders either leveraged existing datasets (Chen et al., 2024) or performed network scanning (van der Ham and van Rijswijk-Deij, 2017) to assess the impact scale of the existing vulnerabilities. The large scale of the notification operations listed in Table 2, unlike disclosure operations in Table 1, has reached significantly higher host numbers ranging from ten to a hundred thousand. In the latest work (Beitis and Vanhoef, [n. d.]), the affected IPv4 hosts have reached a million at an internet-wide scale. On the other hand, the two notification operations (Çetin et al., 2019b; Rodríguez et al., 2021) have fewer affected parties due to the partnership with a medium-sized ISP with its selected customers. In addition, most notifications included more than one vulnerability or security issue of the vulnerable hosts and domains. Hence, the efforts to estimate the exact number of affected parties increased as well.
@@ 5.2.1. Pre-Assessment
Shadowserver Foundation (Beitis and Vanhoef, [n. d.])
@@ 5.2.2. Communication Channel
Selecting a communication channel remains critical (Wogalter et al., 1999) to vulnerability disclosure and notification. Depending on the pre-assessment, a finder can select single to multiple communication channels to reach out to stakeholders at scale, as shown in Table 3. With the current guidelines, there is still no simple decision on channel selection considering the tradeoffs between contact information retrieval and resource capacity, as discussed in all selected works.
@@
For vulnerability disclosure, CVD as best practice is provided in guidelines and gradually adopted by vendors (Walshe and Simpson, 2022b). However, improvements for different scenarios are still being made to best practices. Boucher et al. (Boucher and Anderson, 2022) identified their initial list of affected vendors with bug bounty programs, direct disclosure contacts and outsourced vulnerability platforms. Although they succeeded in most bounty programs, they faced delayed responses or complications in follow-up conversations with outsourced platforms due to their prioritisation of vulnerability reports (Walshe and Simpson, 2022b). They concluded that direct disclosure contacts or bounty programs of vendors might be more effective in extended discussions for their disclosure.
@@
In comparison, Moura et al. (Moura and Heidemann, 2023) experienced late answers and mitigation schedules with a well-known vendor’s direct contact and a bug tracking system. On the other hand, the finders in most work (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Boucher and Anderson, 2022; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]) delegated the notifications to coordinators such as CERT/CC with VINCE and national CSIRTs. Finally, researchers in (Moura and Heidemann, 2023; Van Hove et al., 2023) attended conferences and meetings to reach out to vendors or operators in related communities.
@@
For vulnerability notification, finding an optimised channel to inform the stakeholders at scale turns out to be more challenging than vulnerability disclosure. The most significant difference is that the coordinator delegation to national CSIRTs has been reported as ineffective due to the large scale that went beyond the coverage of the support organisations (Li et al., 2016; Stock et al., 2016a). The finders in work (Çetin et al., 2017a; Stock et al., 2018b, 2016a) exhausted multiple communication channels shown in Table 3 and empirically studied the effectiveness of each channel with the remediation rates by time and feedback from affected parties. From their conclusion, email still remains the prominent channel for conducting large-scale notifications with the best coverage rate despite the obvious pitfalls, such as a high bounce rate, spam filter, and low awareness of recipients.
@@
Similarly, the same concerns were reported in (Durumeric et al., 2014; Li et al., 2016), which used email as their primary channel. In more recent years, notification operations in (Chen et al., 2024; Stivala et al., 2024) adopted best practices derived from the previous work and narrowed their selection to email with a website as support. On the contrary, Maaß et al. (Maass et al., 2021) alternatively selected post and email to compare the effectiveness of both and provided phone support to affected parties. With the considerable amount of manual contact retrieval efforts for email and post, they figured the post could be an effective but costly channel in their nation. Eventually, the researchers in (Chen et al., 2024; Stivala et al., 2024; Maass et al., 2021) all confirmed that the aforementioned issues of notifying stakeholders with email remain.
@@
Across all the selected work, email is the widely used communication channel in initial messages and discussions with vendors and end-users. The main challenges of email are still contact retrieval and low delivery rate. The finders scripted database queries using WHOIS either with purchased database (Stock et al., 2018b, 2016a) or with online query service (Chen et al., 2024), to retrieve email contact information. Although the researchers in (Stock et al., 2018b; Chen et al., 2024) mentioned RDAP as a potential alternative, none attempted it. However, it was reported in (Fernandez et al., 2024) that RDAP has not caught up with the coverage of the existing WHOIS database despite the protocol being introduced to improve contact sharing. Maaß et al. (Maass et al., 2021) confirmed that manual contact retrieval, at best, does not guarantee accurate contact information either. Aside from the contact retrieval issue, email is constantly challenged by high bounce rate (Durumeric et al., 2014; Li et al., 2016), spam filter (Stock et al., 2018b; Chen et al., 2024), and low incentives of recipients to read messages (Çetin et al., 2017a). This eventually results in a low delivery rate to the stakeholders.
@@
Nevertheless, a dedicated scalable notification system may be a solution as researchers and an ISP achieved better results with their walled-garden notifications than email (Çetin et al., 2019b; Rodríguez et al., 2021). Furthermore, the recent ongoing work (Beitis and Vanhoef, [n. d.]) partnered with Shadowserver Foundation, which actively measures the impact and notifies its subscribers at scale. It is a new attempt at a communication channel, and the outcome seems promising.

### v2

@@ 3.1. Stages of Operation
In the literature selection, we cross-reviewed each selected work in detail and aimed to figure out the common procedure for performing vulnerability disclosure and notification. Eventually, we compiled five stages to represent the procedure implemented across the literature. We then used the stages to extract key points, numbers, and remarks from the selected work to better understand efforts, considerations and reviews made in the operations. This results in the following stages to assess the selected publications:
Pre-Assessment – Before a vulnerability disclosure or notification can start, a finder will assess the risk and impact of the discovered vulnerability. We identify the type and number of stakeholders involved and the vulnerabilities, and then extract the impact scale of the vulnerability to understand the preparation required before an operation.
@@
Post-Review – An operation can be reviewed by tracking the remediation rate, feedback from stakeholders, and experiences in each stage. A finder can also reflect on the operation and contribute to best practices. We extract the notified hosts and domains, remediation rates, interactions between the stakeholders, and contributions to best practices of large-scale disclosure and notification operations.
To present an overview of the assessment of each publication, we arrange three tables based on the above five stages. Tables 1 and 2 both contain the publication titles, pre-assessment, and post-review to indicate the efforts and outcomes from each work with different scenarios considered. Table 3 shows the single or multiple communication channels and messaging infrastructure implemented in each reviewed disclosure or notification operation.
The five stages and the three tables serve as our response to the second research question presented in Section 1.1, as well as the foundation to understand the selected publications and extract the insights in the next section.
@@ 4. Lessons Learned
This section presents the insights we have gained from the selected work on large-scale vulnerability disclosure and notification. The main focus is from the finder’s perspective and academic research experience. We also integrate the experience we learn from the communities, support organisations, existing guidelines and standards to present the essence of the disclosure or notification procedure.
With Tables 1, 2, and 3 presented as an overview of disclosure and notification operations, we first discuss the development of best practices from the experiences of the selected work in Section 4.1. Then, we compare how the finders adopted the past best practices and what they learned from their operations in Section 4.2.
@@ 4.1. The Development of Best Practices
Academic security researchers have been seeking optimal methods and strategies as finders based on the disclosure and notification best practices presented in their time, as shown in Table 1 and 2. We have observed changes in their communication channels, messaging infrastructure, contact retrieval, message formulation, and disclosure policies, all of which are influenced by the nature of the vulnerabilities found, notification scalability, and community recommendations.
@@
The essential part of vulnerability disclosure and notification is ensuring messages reach the responsible parties and prompt them to action. As presented in Table 3, the finders used different communication channels and messaging infrastructure to ensure the delivery of the disclosure messages, either on their own or with support organisations, such as national CSIRTs and The Shadowserver Foundation. To maximise notification coverage, all work adopted multiple communication channels, and most of them delegated notifications to support organisations.
@@
The Shadowserver Foundation is a non-profit organisation founded in 2004, driven by the vision of a secure, threat-free internet (Sha, [n. d.]). They do this by scanning the internet for known vulnerabilities, performing notifications, managing sinkholes for malware command-and-control infrastructure, and operating honeypots and honeyclients to monitor threat developments. Over the years, the team at The Shadowserver Foundation has created a large, trusted network with national CSIRTs, ISPs and others to ensure that they are able to reach as many stakeholders as possible. Anyone can sign up to receive threat reports about their domain or IP, provided they can prove ownership. More importantly, The Shadowserver Foundation can support researchers in performing notifications to stakeholders, as described by Beitis and Vanhoef (Beitis and Vanhoef, [n. d.]; The Shadowserver Foundation, [n. d.]).
@@
The Shadowserver Foundation (Beitis and Vanhoef, [n. d.])
@@
The earlier vulnerability notification operations in (Durumeric et al., 2014; Li et al., 2016) focused on email and coordinator delegation notifications, then others in (Çetin et al., 2017a; Stock et al., 2018b, 2016a) extensively implemented diverse communication channels, compared the effectiveness of each empirically, and provided valuable suggestions to best practices on large-scale notifications. These operations served as the foundation for the large-scale notification best practices, especially in the academic security community.
The recent notification operations in (Chen et al., 2024; Stivala et al., 2024) adopted best practices from the prior work. The researchers narrowed down the channel selection and dived into the experiences with email notifications. On the contrary, Maass et al. (Maass et al., 2021) decided to explore alternatives to the best practices with costly manual efforts on email, post, and phone calls. All their results reflected the prior best practices and helped improve notification best practices at scale. In all the above work, locating accurate contact information remains the major challenge regardless of the selected communication channels. On the other hand, finders in (Çetin et al., 2019b; Rodríguez et al., 2021) collaborated with a medium-sized ISP to explore the effectiveness of email and specialised ISP notification systems with different vulnerabilities, which serves as an example in notification with direct intervention from an ISP to their customers.
Vulnerability disclosure to vendors has received promising best practices as implemented in  (Vanhoef and Piessens, 2017a; Boucher and Anderson, 2022; Van Hove et al., 2023). Boucher and Anderson (Boucher and Anderson, 2022) have provided well-structured details on CVE applications, CERT/CC VINCE notification, public press disclosure, and constraints in academic publication. However, despite the stakeholders presenting bug bounty programs and disclosure policies, they encountered slow responses and time constraints in their disclosure operations, as well as in academic publication scheduling. On the other hand, the researchers in the two disclosure operations (Vanhoef and Piessens, 2017a; Van Hove et al., 2023) tackled different software vulnerabilities but encountered the challenge of inviting the same stakeholder to participate in CVD. Last but not least, the two operations in (Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]) cover both vulnerability disclosure and notification. Their experiences revealed the long-lasting challenges of transitioning from disclosure to large-scale notification, as well as the need to explore newer communication channels and strategies to mitigate vulnerabilities in a timely manner.
4.2. The Comparison of Different Scenarios
To understand how the finders in the selected work adopted and contributed to best practices, we describe and compare their experiences based on the operation stages introduced in Section 3.1. The comparison follows the order of the stages: Pre-Assessment, Communication Channel, Messaging Infrastructure, Disclosure Policy and Message, and Post-Review.
4.2.1. Pre-Assessment
All disclosure operations (Vanhoef and Piessens, 2017a; Boucher and Anderson, 2022; Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]) initiated CVD directly to affected vendors with newly found vulnerabilities and later delegated their disclosure to support organisations. Before their disclosure, most finders could identify several vendors with bug bounty (Boucher and Anderson, 2022) or tracking program (Moura and Heidemann, 2023), and a short list of direct disclosure contact information (Vanhoef and Piessens, 2017a; Boucher and Anderson, 2022; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]; Moura and Heidemann, 2023), which ranges from 5 to 14 affected parties due to limited capacity or testing environments. However, most finders faced challenges such as limited contact information of other stakeholders and the potential development to large-scale disclosure. They then consulted national CSIRTs, CERT/CC (VINCE), or NCSC-NL, except for the researchers in (Moura and Heidemann, 2023). The impact scale later increased dramatically from the cooperation with support organisations, which is presented in the Post-Review in Table 1. Despite the lack of contact information and prioritisation in the pre-assessment, the finders could still mitigate the challenges with the support organisations.
The notification operations are vulnerability notifications to the stakeholders, with hosts or domains remaining vulnerable after public disclosure. The finders either leveraged existing datasets (Chen et al., 2024) or performed network scanning (van der Ham and van Rijswijk-Deij, 2017) to assess the impact scale of the existing vulnerabilities or other security issues. The large scale of the notification operations listed in Table 2, unlike disclosure operations in Table 1, has reached significantly higher numbers, ranging from ten to a hundred thousand. In the latest operation in (Beitis and Vanhoef, [n. d.]), the affected IPv4 hosts have reached a million at an internet-wide scale. On the other hand, the two notification operations in (Çetin et al., 2019b; Rodríguez et al., 2021) have fewer affected parties due to the partnership with a medium-sized ISP with its selected customers. In addition, most notifications included more than one vulnerability or security issue of the vulnerable hosts and domains. Hence, the efforts to estimate the exact number of affected parties increased as well.
@@ 4.2.2. Communication Channel
Selecting a communication channel remains critical (Wogalter et al., 1999) to vulnerability disclosure and notification. Depending on the pre-assessment, a finder can select single to multiple communication channels to reach out to stakeholders at scale, as shown in Table 3. With the current guidelines for finders to perform large-scale disclosure or notification, there is still no simple decision on channel selection considering the tradeoffs between contact information retrieval and resource capacity, as discussed in all selected works.
@@
For vulnerability disclosure, CVD as best practice and support mechanisms, such as bug bounty programs, disclosure programs, and disclosure policies, are provided in guidelines and gradually adopted by vendors (Walshe and Simpson, 2022b). However, improvements for different scenarios are still being made to best practices. Boucher and Anderson (Boucher and Anderson, 2022) identified their initial list of affected vendors through bug bounty programs, direct disclosure contacts, and outsourced vulnerability platforms. Although the researchers succeeded in most bounty programs, they encountered delayed responses or complications in follow-up conversations with outsourced platforms due to the latter’s prioritisation of vulnerability reports (Walshe and Simpson, 2022b). They concluded that direct disclosure contacts or bounty programs of vendors might be more effective in extended discussions for their disclosure.
@@
In comparison, Moura and Heidemann (Moura and Heidemann, 2023) experienced late answers and delayed mitigation schedules with a well-known vendor’s direct contact and a bug tracking system. On the other hand, all the finders in (Vanhoef and Piessens, 2017a; Boucher and Anderson, 2022; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]) except for Moura and Heidemann (Moura and Heidemann, 2023) delegated the notifications to coordinators such as CERT/CC with VINCE and national CSIRTs. Finally, finders in (Moura and Heidemann, 2023; Van Hove et al., 2023) attended conferences and meetings to reach out to vendors or operators in related communities.
@@
For vulnerability notification, finding an optimised channel to inform the stakeholders at scale turns out to be more challenging than vulnerability disclosure. The most significant difference is that the coordinator delegation to national CSIRTs has been reported as ineffective due to the large scale that went beyond the coverage of the support organisations, as reported in (Li et al., 2016; Stock et al., 2016a). The finders in (Çetin et al., 2017a; Stock et al., 2018b, 2016a) exhausted multiple communication channels shown in Table 3 and empirically studied the effectiveness of each channel with the remediation rates by time and feedback from affected parties. In the three publications, they all concluded that email still remains the prominent channel for conducting large-scale notifications with the best coverage rate despite the obvious pitfalls, such as a high bounce rate, spam filter, and low awareness of recipients.
@@
Similarly, the same concerns were reported by finders in (Durumeric et al., 2014; Li et al., 2016), which used email as their primary channel. In recent years, finders in (Chen et al., 2024; Stivala et al., 2024) adopted best practices derived from the previous work and narrowed their selection to email with a website as support. On the contrary, Maass et al. (Maass et al., 2021) alternatively selected post and email to compare the effectiveness of both and provided phone support to notified parties. Given the considerable manual effort required for email and post contract information retrieval, they concluded that the post could be an effective but costly channel in their nation. Eventually, the finders in (Chen et al., 2024; Stivala et al., 2024; Maass et al., 2021) all confirmed that the aforementioned issues of notifying stakeholders via email remain.
@@
Across all the selected publications, email is the widely used communication channel in initial messages and discussions with vendors and end-users. The main challenges of email are still contact retrieval and low delivery rate. The finders scripted database queries using WHOIS either with a purchased database as in (Stock et al., 2018b, 2016a) or with an online query service as in (Chen et al., 2024), to retrieve email contact information. Although the finders in (Stock et al., 2018b; Chen et al., 2024) mentioned RDAP as a potential alternative, none attempted it. However, Fernandez et al.  (Fernandez et al., 2024) reported that RDAP has not caught up with the coverage of the existing WHOIS database despite the protocol being introduced to improve contact sharing. Moreover, Maass et al. (Maass et al., 2021) confirmed that manual contact retrieval, at best, does not guarantee accurate contact information either. Aside from the contact retrieval issue, email is constantly challenged by high bounce rate (Durumeric et al., 2014; Li et al., 2016), spam filter (Stock et al., 2018b; Chen et al., 2024), and low incentives of recipients to read messages (Çetin et al., 2017a). This eventually results in a low delivery rate to the stakeholders.
@@
Nevertheless, a dedicated scalable notification system with a notification partner can be a solution. Finders in  (Çetin et al., 2019b; Rodríguez et al., 2021) teamed up with an ISP and achieved better results with the ISP’s walled-garden notifications than email. On the other hand, Beitis and Vanhoef (Beitis and Vanhoef, [n. d.]) partnered with The Shadowserver Foundation, which actively measures the mitigation progress and notifies its subscribers at scale. Their notification is a new attempt at a communication channel, and the outcome seems promising.